OpenStudy (anonymous):

A circle has a diameter of 34 yd . What is its area? Use 3.14 for (I DON'T KNOW THIS SYMBOL:(!!! OMG) choices: 106.76 sq yd 213.52 sq yd 453.73 sq yd 907.46 sq yd

OpenStudy (anonymous):

THIS SYMBOL \[pi\]

OpenStudy (lgbasallote):

you mean \(\pi\)?

OpenStudy (anonymous):

Yes.

OpenStudy (lgbasallote):

area = \(\pi (\frac{d}{2})^2\)

OpenStudy (lgbasallote):

where d = diameter...\(\pi = 3.14\) input them in the calculator

OpenStudy (lgbasallote):

\(r^2 \ne diameter\)

OpenStudy (anonymous):

106.76 sq yd 213.52 sq yd 453.73 sq yd 907.46 sq yd

OpenStudy (lgbasallote):

ugh just put it in the calculator =_=

OpenStudy (anonymous):

i'm out of this site... peace.. it's lame!